H8 Q6 Consider a competitive market served by many domestic and foreign firms. The domestic demand...

H8 Q6

Consider a competitive market served by many domestic and foreign firms. The domestic demand for these firms’ product is Qd = 1400 - 2.5P. The supply function of the domestic firms is QSD = 200 + 1.5P, while that of the foreign firms is QSF = 200

Instructions: Round your answers for equilibrium price to the nearest penny (two decimal places). Round your answers for equilibrium quantity to one decimal place.

a. Determine the equilibrium price and quantity under free trade.

Equilibrium price: $
Equilibrium quantity: units


b. Determine the equilibrium price and quantity when foreign firms are constrained by a 100-unit quota.

Equilibrium price: $
Equilibrium quantity: units


c. Are domestic consumers better or worse off as a result of the quota?

Which one?

Worse off

Better off

Neither better nor worse off




d. Are domestic producers better or worse off as a result of the quota?

Which one?

Better off

Worse off

Neither better nor worse off
Answer #1

(a) Under free trade, Qd = QSD + QSF

1400 - 2.5P = 200 + 1.5P + 200

1400 - 2.5P = 400 + 1.5P

1400 - 400 = 1.5P + 2.5P

1000 = 4P

P = 1000 / 4

P = 250 (Equilibrium Price)

Q = 1400 - (2.5 x 250)

Q= 1400 - 625

Q= 775 (Equilibrium Quantity)

(b) QSF = 100 (Foreign firms are constrained by 100 unit quota)

1400 - 2.5P = 200 + 1.5P + 100

1400 - 2.5P = 300 + 1.5P

1400 - 300 = 1.5P + 2.5P

1100 = 4P

P = 1100 / 4

P = 275 (Equilibrium price)

Q = 1400 - (2.5 x 275)

Q= 1400 - 687.5

Q= 712.5 (Equilibrium Quantity)

(c) Since price has increased due to quota, domestic consumers are worse off (since consumer surplus will decrease).

(d) Since price has increased due to quota, domestic producers are better off (since producer surplus will increase).